OpenStudy (anonymous):

Factor the trinomial: 2x2 + 7x + 3 A. (2x - 1)(2x + 3) B. (2x + 3)(x + 1) C. (2x + 1)(x + 3) D. (2x + 1)(2x + 3)

OpenStudy (anonymous):

C

OpenStudy (jadeishere):

Explain. You can't just give the answer like that There's two ways you could figure this out; 1) solve each of the factors 2) factor the trinomial normally

OpenStudy (anonymous):

First one: find two numbers which add up to 7 and multiply to make 6. Those are 6 and 1 Replace the middle term by those numbers: 2x^2 + (6 + 1)x + 3 = 0 2x^2 + 6x + x + 3 = 0 Now factor the two groups: 2x(x+3) + (x+3) = 0 Your factored equation is: (2x + 1)(x + 3) = 0 with solutions x=-1/2 and x=-3 For the second: Move the right hand side to the left and simplify. You'll get: 6d^2 + 11d - 35 = 0 Now we need 2 numbers which add up to 11 and multiply to make -210. Those are 21 and -10. Now do the same steps as before: 6d^2 + (21 - 10)d - 35 = 0 6d^2 + 21d - 10d - 35 = 0 3d(2d + 7) - 5(2d + 7) = 0 Final equation: (3d - 5)(2d + 7) = 0 Solutions: d = 5/3 or -7/2

OpenStudy (anonymous):

(2x + 1)(x + 3) = 2x^2 + 6x +1x +3 =2x^2 + 7x + 3 DONE.